OpenStudy (anonymous):

What is the inverse function of f^(-1(x) ) if f(x)=√((3x-1) )+2

OpenStudy (jhannybean):

To find inverse functions, 1. set f(x) = y = \(\sqrt{3x-1} +2\) 2. switch the x and y around : x = \(\sqrt{3y-1}+2\) 3. solve for y now

OpenStudy (anonymous):

can you walk me through it please

OpenStudy (jhannybean):

\[x -2 = \sqrt{3y-1}\]\[(x-2)^2 = (\sqrt{3y-1})^2\]\[(x-2)^2 = 3y -1\]\[(x-2)^2 +1 =3y\]\[y= \frac{(x-2)^2+1}{3}\]

OpenStudy (jhannybean):

Which part are you confused about currently? :o

OpenStudy (anonymous):

wow. thanks.....I forgot about squaring both sides...I was only squaring one side

OpenStudy (jhannybean):

Yeah, at first I thought hey, we got it in terms of x = something... why not just square both sides?

OpenStudy (jhannybean):

then i realized that if i squared the right side with the 2 still there, it would be really messy and complicated, so i subtracted 2 from both sides before squaring both sides and getting rid of the square root.

OpenStudy (anonymous):

yeah! thank!(:

OpenStudy (jhannybean):

First step : change f(x) to y 2nd step: switch the x and y to indicate you're taking the inverse of the function 3rd step: solve for y after inverting the two variables 4th step: once you get your equation in terms of y (again) change it to \(f^{-1} (x)\) to indicate that it is the inverse.

OpenStudy (anonymous):

Can you walk me through the steps on finding the domain and rang for f(x)=logbase4(x+2)?

OpenStudy (anonymous):

*range